OpenStudy (eriyalovesu):

MEDAL The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 0.69(1.03)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 10,100 10,201 10,303.01 10,406.04

OpenStudy (eriyalovesu):

Which product recorded a greater percentage change in price over the previous year? Justify your answer. is Part B

OpenStudy (eriyalovesu):

product A is decreasing by 0.63 right?

OpenStudy (eriyalovesu):

What about part b?

OpenStudy (eriyalovesu):

that doesn't help with part b tho

OpenStudy (allen27):

for B, take the price of year 2 - the price of year 1, you get 10,201 - 10,100=101 then take the price of year 3- the price of year 2, you get 10,303.01-10201=102.01 (a little bit up, right?) do the same with year 4, 3 10,406.04-10,303.01=103.03 (a little bit up again, right?)

OpenStudy (allen27):

@phi

OpenStudy (phi):

First, is the equation for part A \[ f(x) = 0.69(1.03)^x\] where x is an exponent? if so , then f(x) is growing by 3%

OpenStudy (phi):

for part B, you first find the difference (see allen's work) then divide by the original amount \[ \frac{101}{10,100} \] then change that to a percent by multiplying by 100 (and adding % sign)

OpenStudy (eriyalovesu):

okay

OpenStudy (eriyalovesu):

and its not his work he copied and pasted it all of it from another person